LHC 2010 - latest news

Week 19: Machine Coordinators:    Mike Lamont - Ralph Assmann

Latest plan: link PDF link XLS (version of 14 May, 13:30)

List of End of Fill studies (in work)

List of MD requests (in work)

Day Short summary/plan Link to reports from 8:30 and 17:00 meetings
Monday 17th May 2010
  • 00h07: Recovery and injection setup: 6x2e10 per beam.
  • 06h00: Stable beams (6 x 2e10 / beam)
Report from 8h30 meeting
Sunday 16th May 2010
  • 14h00: End of stable beams (6x2e10/beam). End of fill studies: Single beam for CMS van der Meer scan. Loss map.
  • 15h00: Recovery and injection setup: 1x1e10 per beam.
  • 17h49: Start test ramp with low intensity and "virgin" orbit and tunes. Dumped with TCDQ position interlocks during ramp.
  • 19h00: Recovery and injection setup: 1 x 1e11.
  • 21h00: Start of high intensity studies at 450 GeV: chromaticity, Landau octupoles
  • 23h06: Start ramp with 1 beam of 1.2e11 in 1 bunch. Octupoles used for Landau damping.
  • 23h50: 1.2e11 bunch at 3.5 TeV without losses.
  • 00h00: Zeroing octupoles leads to beam instability, beam losses and beam dump.
Report from 9h00 meeting
Saturday 15th May 2010
  • 02h20: Recovery and injection setup: 4 x 2e10.
  • 06h03: Start of ramp for stable beams.
  • 09h00: Injection of nominal bunches (1e11 per beam).
  • 09h48: Start of ramp. 60% of beam survived to 3.5 TeV. Strong beam losses and excitation of synchrotron sidebands during ramp for both beams.
  • 12h00: Recovery and injection setup: 6 x 2e10.
  • 16h54: Stable beams.
Report from 9h00 meeting
Friday 14th May 2010
  • 2h00: Prepare fill for stable beams (2x2e10 per beam). 1 beam dumped on ramp.
  • 8h00: Loading new SIS software as final step for increasing intensity. Prepare fill with higher intensity.
  • 13h00: Start stable beams with factor 2 higher intensities (4x2e10 per beam).
  • 23h41: Beams dumped. RF cavity tripped due to quench.
Report from 9h00 meeting
Thursday 13th May 2010
  • 01h40: Start of ramp for beta beat studies. 1 beam lost.
  • 02h47: Requested beam dump at 3.5 TeV. Ramp down and cycle.
  • 05h47: Start of ramp for beta beat studies.
  • 07h20: Start of RF studies at 3.5 TeV.
  • 09h00: Beam dump and recovery.
  • 12h00: Injection protection setup for higher intensities.
  • 14h30: Collisions @450 GeV (no stable beams) to ATLAS for calibration of liquid argon calorimeter.
  • 15h00: Orbit feedback checks. Collimator energy-dependent gap thresholds.
  • 18h00: Ramp with collimators and feedback (low intensity).
  • 22h07: Lost cryo mantain and start in MSR6.
  • 23h45: BPM sensitivity checks to intensity at 450 GeV.
Report from 9h00 meeting
Wednesday 12th May 2010
  • 01h30: Injection and setup with beam.
  • 04h20: Start of study of BPM intensity and bunch length dependence.
  • 05h30: PS problem. No beam for LHC.
  • 08h00: PS back. Injection and setup with beam.
  • 09h00: LBDS Test. Injection protection check.
  • 10h00: Check of new MQW functions with beam. RF synchronization test for LBDS. Collimation ramp without beam.
  • 17h00: Ramp with collimators closing and orbit feedback.
  • 19:00: Rampdown and cycle.
  • 21h00: Beam setup at injection.

Report from 8h30 meeting

Report from LMC

Tuesday 11th May 2010
  • 17h00: Accesses in shadow of cryo recovery finished.
  • 17h30: Pre-cycle started.
  • 19h00: Preparing access for intervention of RQTL11.R2B2.
Report from the 8h30 meeting
Monday 10th May 2010
  • 11h00: Beams dumped by LBDS trigger. End of 3.5 TeV physics.
  • 11h00-14h00: Recovery and prepare for beam dump studies at 450 GeV.
  • 14h00-23h00: Beam dump studies at injection.
  • 23h00: Cold compressor problem in 18.
 Report from the 8h30 meeting